Miguel works three times faster than Steven together they do a job in nine hours How long does it take Steven working alone to do the same job
Found 2 solutions by checkley79, Alan3354:
Answer by checkley79(3341)

s=x
m=3x
1/x+1/3x=1/9
(3+1)/3x=1/9
4/3x=1/9
3x=4*9
3x=36
x=12 hours for Miguel working alone.
m=3*12=36 hours for Steven working alone.
Proof:
1/12+1/36=1/9
(3+1)/36=1/9
4/36=1/9
1/9=1/9

Answer by Alan3354(69443)

Miguel works three times faster than Steven together they do a job in nine hours How long does it take Steven working alone to do the same job
----------------------
3 times faster = 4 times as fast
---
1/M + 1/S = 1/9
1/4S + 1/S = 1/9
5/4S = 1/9
S = 45/4 hours for Steve alone
================================
3 times faster is an increase of 300%
An increase of 100% is 1 times faster, 2x as fast
An increase of 200% is 2 times faster, 3x as fast
An increase of3100% is 3 times faster, 4x as fast
-----
"times faster" is not the same as "times as fast"